Suppose the demand equation for shale gas is Qd = 10P ^ −1.8 and the supply equation Qs=2P ^ 0.2
What is the equilibrium price and equilibrium quantity?
1
Expert's answer
2020-11-16T06:53:07-0500
"Q_D=10p^{-1.8}"


"Q_S=2p^{0.2}"


"10p^{-1.8}=2p^{0.2}"


"p^2=5"


"p=\sqrt 5"

"Q=\sqrt [10] {5}"